# Complete NLO BFKL impact factors for quarkonium hadroproduction in NRQCD: the case of $^{1}S_{0}^{[1]}$, $^{1}S_{0}^{[8]}$, and $^{3}S_{1}^{[8]}$ states

### v2

@@ 1 Introduction
The current frontier is the NLL precision [^2] Some pieces of the next-to-NLO kernel have been recently found in $\mathcal{N}=4$ SYM Byrne:2022wzk , in pure-gauge QCD DelDuca:2021vjq and in full QCD Caola:2021izf ; Falcioni:2021dgr ; Fadin:2023roz ; Abreu:2024mpk ; Buccioni:2024gzo ; Abreu:2024xoh ., which requires the knowledge at Next-to-Leading Order (NLO) corrections to both the kernel of the BFKL equation  Fadin:1998py ; Ciafaloni:1998gs ; Fadin:1998jv ; Fadin:2000kx ; Fadin:2000hu ; Fadin:2004zq ; Fadin:2005zj and to the impact factors (IFs) Fadin:1999de ; Fadin:1999df ; Ciafaloni:1998kx ; Ciafaloni:1998hu ; Ciafaloni:2000sq ; Bartels:2001ge ; Bartels:2002yj ; Caporale:2011cc ; Ivanov:2012ms ; Colferai:2015zfa ; Chirilli:2012jd ; Hentschinski:2014bra ; Bartels:2000gt ; Bartels:2001mv ; Bartels:2002uz ; Bartels:2003zi ; Bartels:2004bi ; Fadin:2001ap ; Ivanov:2012iv ; Nefedov:2024swu ; Nefedov:2019mrg ; Hentschinski:2020tbi ; Celiberto:2022fgx ; Fucilla:2024cpf ; Celiberto:2024bbv . Currently, there are only a few processes for which the full NLL BFKL cross section has been computed, such as the production of Mueller-Navelet MN:87 ; Colferai:2010wu ; Caporale:2012ih ; Ducloue:2013hia ; Ducloue:2013bva ; Ducloue:2014koa ; Caporale:2014gpa ; Ducloue:2015jba ; Mueller:2015ael or Mueller-Tang Mueller:1992pe ; Colferai:2023hje dijets. Those inclusive processes to be studied at the LHC all involve large transverse momenta, typically above tens of GeV, which reduces the size of BFKL effects, such as azimuthal de-correlation. In the case of processes involving quarkonia, $\cal Q$}, it is important to note that existing NLL cross sections are computed Celiberto:2022dyf ; Celiberto:2025euy at leading-power in $P_{T}$, that is neglecting corrections scaling like to $M_{\cal Q}^{2}/P_{T}^{2}$. This is also known as the fragmentation approximation.
@@
In the present paper, we wish to demonstrate that it is possible to compute quarkonium-related cross sections up to NLL without this restriction when the quarkonium is produced at large forward (backward) rapidities in association with another particles produced at backward (forward) rapidities, in the exact analogy with the Mueller-Navelet dijets. To do so, we rely on the Non-Relativistic QCD (NRQCD) factorisation framework Bodwin:1994jh which has become a standard formalism for the description of hadronisation of heavy quark pair ($Q\bar{Q}$) into heavy quarkonium ($\cal Q$}Brambilla:2004wf ; Brambilla:2010cs ; Lansberg:2019adr ; OniumEIC and which is based on a double expansion on the strong coupling constant $\alpha_{S}$ and the relative velocity of the heavy quarks in the quarkonium rest frame, $v$. Within NRQCD, the $Q\bar{Q}$-pair is produced at short distances in various states $Q\bar{Q}[m]$, denoted $m={}^{2S+1}L^{[1,8]}_{J}$ using the spectroscopic notation with the total spin $S$, orbital momentum $L$, total angular momentum $J$ and with colour-singlet ([1]) or colour-octet ([8]) quantum numbers. The production of the state $m$ is described perturbatively, while the probability of hadronisation of the state $m$ into quarkonium $\cal Q$} is given by the Long-Distance Matrix Elements (LDMEs) of NRQCD: $\langle{\cal O}^{\cal Q}[m]\rangle$. The velocity-scaling rules for LDMEs contributing to production of various physical quarkonium states are summarised in tab. 1 together with the $\alpha_{s}$ orders of the corresponding LO IFs.
@@
The goal of the present paper is to compute the NLO correction to the IF $V^{(m,1)}_{a}$ for $a=g,q$ and $m={}^{1}S_{0}^{[1]}$, $^{1}S_{0}^{[8]}$ and $^{3}S_{1}^{[8]}$. It turns out that the key technical differences in the computation concern the distinction between CS and CO cases, while the difference between $S=0$ and $S=1$ cases is marginal. Therefore in the next two sections we present respectively the results for the NLO IFs of production of the CS and CO $Q\bar{Q}$-state and then proceed with explaining the details of their derivation.
@@ 3.1 The subtraction approach for the real-emission part
The IF is a function of $z$, ${\bm{q}}_{T}$ and ${\bm{p}}_{T}$, which should be integrated over $z$ with the PDFs in eqn. (6) and over ${\bm{q}}_{T}$ with the Green’s function in the eqn. (7). However these integrations will contain collinear (${\bm{p}}_{T}-{\bm{q}}_{T}\to 0$), soft (${\bm{p}}_{T}-{\bm{q}}_{T}\to 0$ and $z\to 1$) and rapidity ($z\to 1$) divergences which should be isolated. These will either cancel against the corresponding $1/\epsilon$ and $1/\epsilon^{2}$ poles in the virtual corrections to the IF, computed in Ref. Nefedov:2024swu , or should be properly subtracted according to the chosen factorisation scheme. The real-emission contributions to the IF are too complicated for this purpose. Therefore in the present paper we use the subtractive approach, similar to one used in Ref. Hentschinski:2020tbi to isolate the divergences. We add and subtract a suitable function, $\cal J$}, which removes all singularities from the reduced squared matrix element, $\tilde{H}$, and for which the divergences can be computed in closed form:
@@ 5.1 Extracting divergences from the CS subtraction term
$V_{a\text{, ST}}^{(m,1)}$ $=$ $\frac{2z|{\bm{p}}_{T}|\mu^{2\epsilon}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\int d^{2-2\epsilon}{\bm{k}}_{T}\,{\cal J}^{(m)}_{Ra}({\bm{k}}_{T},{\bm{p}}_{T},z)\delta^{(2)}({\bm{q}}_{T}-{\bm{k}}_{T}-{\bm{p}}_{T}).$ (39)
@@
$V_{a\text{, ST-2}}^{(m,1)}$ $=$ $\frac{2z|{\bm{p}}_{T}|\mu^{2\epsilon}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\int d^{2-2\epsilon}{\bm{k}}_{T}\frac{{\bm{p}}_{T}^{2}{\cal J}^{(m)}_{Ra}({\bm{k}}_{T},{\bm{p}}_{T},z)}{{\bm{p}}_{T}^{2}+{\bm{k}}_{T}^{2}}.$ (42)
@@
The appearance of $K_{\text{BFKL}}$ in this equation is a consequence of our computation procedure and is not related to any evolution.
@@
$\hskip-28.45274ptV_{g\text{, ST-2}}^{(m,1)}$ $=$ $C_{\overline{MS}}z|{\bm{p}}_{T}|\ h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})$
@@
$\hskip-28.45274ptV_{q\text{, ST-2}}^{(m,1)}$ $=$ $C_{\overline{MS}}z|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\int\limits_{0}^{\infty}\frac{d{\bm{k}}_{T}^{2}}{{\bm{k}}_{T}^{2}}\bigg(\frac{{\bm{k}}_{T}^{2}}{\mu^{2}}\bigg)^{-\epsilon}\frac{{\bm{p}}_{T}^{2}}{{\bm{p}}_{T}^{2}+{\bm{k}}_{T}^{2}}\left[P_{gq}(z)-\epsilon C_{F}z\right],\,$ (49)
@@
$V_{g\text{, ST-2}}^{(m,1)}=C_{\overline{MS}}z|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\,$ (50)
@@
$\hskip-28.45274ptV_{q\text{, ST-2}}^{(m,1)}=C_{\overline{MS}}z|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\bigg(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\bigg)^{\epsilon}\bigg(-\frac{1}{\epsilon}+O(\epsilon)\bigg)\left[P_{gq}(z)-\epsilon C_{F}z\right],$ (51)
@@ 5.2 Loop correction and renormalisation
$V_{g\text{, V}}^{({}^{1}S_{0}^{[1]},1)}$ $=C_{\overline{MS}}\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{{}^{1}S_{0}^{[1]}}^{(0)}({\bm{p}}_{T}^{2})\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}$
@@
where $F_{{}^{1}S_{0}^{[1]}}(\tau)$ is the finite part of the one-loop correction, see eqn. (27). In Ref. Nefedov:2024swu , the one-loop amplitude was properly LSZ-amputated with respect to the on-shell particles, so one needs to add to it just two UV counterterms:
@@
$C_{\overline{MS}}\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\left\{-\frac{1}{\epsilon}\left[\beta_{0}+3C_{F}\right]-2C_{F}\left(2+\frac{3}{2}\ln\frac{\mu_{R}^{2}}{m_{Q}^{2}}\right)\right\},$ (54)
@@
$\hskip-14.22636ptV_{g\text{, VR}}^{({}^{1}S_{0}^{[1]},1)}=C_{\overline{MS}}\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{{}^{1}S_{0}^{[1]}}^{(0)}({\bm{p}}_{T}^{2})\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}$ (55)
@@
Note that the terms underlined by a single line cancel against the corresponding terms in eqn. (50). The remaining divergences, underlined by two lines in eqns. (50) and (5.2), will be cancelled by the counterterm corresponding to the transition to the BFKL rapidity-factorisation scheme, discussed in the next subsection.
@@ 5.3 BFKL counterterm for the virtual part
$V_{g\text{, BFKL-V}}^{(m,1)}=C_{\overline{MS}}\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}$
@@
As anticipated above, the doubly-underlined divergent terms in eqn. (57) cancel the corresponding terms coming from eqns. (50) and (5.2). These terms in eqn. (57) have their origin from the non-rapidity-divergent part of the one-loop correction to the reggeised gluon propagator in the EFT Chachamis:2012cc ; Nefedov:2019mrg . The $\ln r/\epsilon$ term, highlighted with the wavy line in eqn. (57) will cancel the IR divergence of the scheme-transformation term for the real correction, derived in the next subsection (see eqn. (69)).

### v3

@@ 1 Introduction
The current frontier is the NLL precision [^2] Some pieces of the next-to-NLO kernel have been recently found in $\mathcal{N}=4$ SYM Byrne:2022wzk , in pure-gauge QCD DelDuca:2021vjq and in full QCD Caola:2021izf ; Falcioni:2021dgr ; Fadin:2023roz ; Abreu:2024mpk ; Buccioni:2024gzo ; Abreu:2024xoh ., which requires the knowledge at Next-to-Leading Order (NLO) corrections to both the kernel of the BFKL equation  Fadin:1998py ; Ciafaloni:1998gs ; Fadin:1998jv ; Fadin:2000kx ; Fadin:2000hu ; Fadin:2004zq ; Fadin:2005zj and to the impact factors (IFs) Fadin:1999de ; Fadin:1999df ; Ciafaloni:1998kx ; Ciafaloni:1998hu ; Ciafaloni:2000sq ; Bartels:2001ge ; Bartels:2002yj ; Caporale:2011cc ; Ivanov:2012ms ; Colferai:2015zfa ; Chirilli:2012jd ; Hentschinski:2014bra ; Bartels:2000gt ; Bartels:2001mv ; Bartels:2002uz ; Bartels:2003zi ; Bartels:2004bi ; Fadin:2001ap ; Ivanov:2012iv ; Nefedov:2024swu ; Nefedov:2019mrg ; Hentschinski:2020tbi ; Celiberto:2022fgx ; Fucilla:2024cpf ; Celiberto:2024bbv . Currently, there are only a few processes for which the full NLL BFKL cross section has been computed, such as the production of Mueller-Navelet MN:87 ; Colferai:2010wu ; Caporale:2012ih ; Ducloue:2013hia ; Ducloue:2013bva ; Ducloue:2014koa ; Caporale:2014gpa ; Ducloue:2015jba ; Mueller:2015ael or Mueller-Tang Mueller:1992pe ; Colferai:2023hje dijets. Those inclusive processes to be studied at the LHC all involve large transverse momenta, typically above tens of GeV, which reduces the size of BFKL effects, such as azimuthal de-correlation. In the case of processes involving quarkonia, $\cal Q$}, it is important to note that existing NLL cross sections are computed Celiberto:2022dyf ; Celiberto:2025euy ; Celiberto:2023fzz at leading-power in $P_{T}$, that is neglecting corrections scaling like to $M_{\cal Q}^{2}/P_{T}^{2}$. This is also known as the fragmentation approximation.
@@
In the present paper, we wish to demonstrate that it is possible to compute quarkonium-related cross sections up to NLL without this restriction when the quarkonium is produced at large forward (backward) rapidities in association with another particles produced at backward (forward) rapidities, in the exact analogy with the Mueller-Navelet dijets. To do so, we rely on the Non-Relativistic QCD (NRQCD) factorisation framework Bodwin:1994jh which has become a standard formalism for the description of hadronisation of heavy quark pair ($Q\bar{Q}$) into heavy quarkonium ($\cal Q$}Brambilla:2004wf ; Brambilla:2010cs ; Lansberg:2019adr ; OniumEIC and which is based on a double expansion on the strong coupling constant $\alpha_{S}$ and the relative velocity of the heavy quarks in the quarkonium rest frame, $v$. Within NRQCD, the $Q\bar{Q}$-pair is produced at short distances in various states $Q\bar{Q}[m]$, denoted $m={}^{2S+1}L^{[1,8]}_{J}$ using the spectroscopic notation with the total spin $S$, orbital momentum $L$, total angular momentum $J$ and with colour-singlet (CS,[1]) or colour-octet (CO,[8]) quantum numbers. The production of the state $m$ is described perturbatively, while the probability of hadronisation of the state $m$ into quarkonium $\cal Q$} is given by the Long-Distance Matrix Elements (LDMEs) of NRQCD: $\langle{\cal O}^{\cal Q}[m]\rangle$. The velocity-scaling rules for LDMEs contributing to production of various physical quarkonium states are summarised in tab. 1 together with the $\alpha_{s}$ orders of the corresponding LO IFs.
@@
The goal of the present paper is to compute the NLO correction to the IF $V^{(m,1)}_{a}$ for $a=g,q$ and $m={}^{1}S_{0}^{[1]}$, $^{1}S_{0}^{[8]}$ and $^{3}S_{1}^{[8]}$. It turns out that the key technical differences in the computation concern the distinction between colour-singlet(CS, [1]) and colour-octet (CO, [8]) cases, while the difference between $S=0$ and $S=1$ cases is marginal. Therefore in the next two sections we present respectively the results for the NLO IFs of production of the CS and CO $Q\bar{Q}$-state and then proceed with explaining the details of their derivation.
@@ 3.1 The subtraction approach for the real-emission part
The IF is a function of $z$, ${\bm{q}}_{T}$ and ${\bm{p}}_{T}$, which should be integrated over ${\bm{q}}_{T}$ with the Green’s function in the eqn. (7) and over $x_{1,2}$ with the PDFs in eqn. (6). We recall (eqn. (5)) that the variables $x_{1,2}$ are related to the variables $z_{1,2}=M_{T1,2}e^{y_{1,2}}/(x_{1,2}\sqrt{S})$. These integrations will contain collinear (${\bm{p}}_{T}-{\bm{q}}_{T}\to 0$), soft (${\bm{p}}_{T}-{\bm{q}}_{T}\to 0$ and $z\to 1$) and rapidity ($z\to 1$) divergences which should be isolated. These will either cancel against the corresponding $1/\epsilon$ and $1/\epsilon^{2}$ poles in the virtual corrections to the IF, computed in Ref. Nefedov:2024swu , or should be properly subtracted according to the chosen factorisation scheme. The real-emission contributions to the IF are too complicated for this purpose. Therefore in the present paper we use the subtractive approach, similar to one used in Ref. Hentschinski:2020tbi to isolate the divergences. We add and subtract a suitable function, $\cal J$}, which removes all singularities from the reduced squared matrix element, $\tilde{H}$, and for which the divergences can be computed in closed form:
@@ 5.1 Extracting divergences from the CS subtraction term
$V_{a\text{, ST}}^{(m,1)}$ $=$ $\frac{2z|{\bm{p}}_{T}|{\mu^{2\epsilon}}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\int d^{2-2\epsilon}{\bm{k}}_{T}\,{\cal J}^{(m)}_{Ra}({\bm{k}}_{T},{\bm{p}}_{T},z)\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{k}}_{T}-{\bm{p}}_{T}).$ (39)
@@
$V_{a\text{, ST-2}}^{(m,1)}$ $=$ $\frac{2z|{\bm{p}}_{T}|{\mu^{2\epsilon}}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\int d^{2-2\epsilon}{\bm{k}}_{T}\frac{{\bm{p}}_{T}^{2}{\cal J}^{(m)}_{Ra}({\bm{k}}_{T},{\bm{p}}_{T},z)}{{\bm{p}}_{T}^{2}+{\bm{k}}_{T}^{2}}.$ (42)
@@
The appearance of $K_{\text{BFKL}}$ in this equation may be surprising since $K_{\text{BFKL}}$ is the coefficient of the rapidity divergence which is absent in the quark case. However, in our case it appears just as an artifact of our procedure of isolating the collinear singularity at $\bm{k}_{T}\to 0$.
@@
$\hskip-28.45274ptV_{g\text{, ST-2}}^{(m,1)}$ $=$ $C_{\overline{MS}}z|{\bm{p}}_{T}|\ h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})$
@@
$\hskip-28.45274ptV_{q\text{, ST-2}}^{(m,1)}$ $=$ $C_{\overline{MS}}z|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\int\limits_{0}^{\infty}\frac{d{\bm{k}}_{T}^{2}}{{\bm{k}}_{T}^{2}}\bigg(\frac{{\bm{k}}_{T}^{2}}{\mu^{2}}\bigg)^{-\epsilon}\frac{{\bm{p}}_{T}^{2}}{{\bm{p}}_{T}^{2}+{\bm{k}}_{T}^{2}}\left[P_{gq}(z)-\epsilon C_{F}z\right],\,$ (49)
@@
$V_{g\text{, ST-2}}^{(m,1)}=C_{\overline{MS}}z|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\,$ (50)
@@
$\hskip-28.45274ptV_{q\text{, ST-2}}^{(m,1)}=C_{\overline{MS}}z|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})$ (51)
$\times\bigg(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\bigg)^{\epsilon}\bigg(-\frac{1}{\epsilon}+O(\epsilon)\bigg)\left[P_{gq}(z)-\epsilon C_{F}z\right],$
@@ 5.2 Loop correction and renormalisation
$V_{g\text{, V}}^{({}^{1}S_{0}^{[1]},1)}$ $=C_{\overline{MS}}\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{{}^{1}S_{0}^{[1]}}^{(0)}({\bm{p}}_{T}^{2})\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}$
@@
where $F_{{}^{1}S_{0}^{[1]}}^{\text{(virt.)}}(\tau)$ is the finite part of the one-loop correction, see eqn. (27). In Ref. Nefedov:2024swu , the one-loop amplitude was properly LSZ-amputated with respect to the on-shell particles, so one needs to add to it just two UV counterterms:
@@
$C_{\overline{MS}}\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\left\{-\frac{1}{\epsilon}\left[\beta_{0}+3C_{F}\right]-2C_{F}\left(2+\frac{3}{2}\ln\frac{\mu_{R}^{2}}{m_{Q}^{2}}\right)\right\},$ (54)
@@
$\hskip-14.22636ptV_{g\text{, VR}}^{({}^{1}S_{0}^{[1]},1)}=C_{\overline{MS}}\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{{}^{1}S_{0}^{[1]}}^{(0)}({\bm{p}}_{T}^{2})\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}$ (55)
@@
Note that the terms underlined by a single line cancel against the corresponding terms in eqn. (50). The remaining divergences, underlined by two lines in eqns. (50) and (55), will be cancelled by the counterterm corresponding to the transition to the BFKL rapidity-factorisation scheme, discussed in the next subsection.
@@ 5.3 BFKL counterterm for the virtual part
$V_{g\text{, BFKL-V}}^{(m,1)}=C_{\overline{MS}}\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}$
@@
As anticipated above, the doubly-underlined divergent terms in eqn. (57) cancel the corresponding terms coming from eqns. (50) and (55). These terms in eqn. (57) have their origin from the non-rapidity-divergent part of the one-loop correction to the reggeised gluon propagator in the EFT Chachamis:2012cc ; Nefedov:2019mrg . The $\ln r/\epsilon$ term, highlighted with the wavy line in eqn. (57) will cancel the IR divergence of the scheme-transformation term for the real correction, derived in the next subsection (see eqn. (69)).